it is now the beginning of a year. An investor is considering the purchase of a 7 percent (coupon rate), 10-year bond that is presently priced to yield 12 percent (i.e. market interest rate is 12 percent). Based on extensive analysis of market interest rates, he thinks rates will fall so that at the end of this year the market yield of this issue will drop to 8 percent. If his expectations are correct, what kind of realized return will the investor earn by purchasing the bond today and selling the bond at the end of the year? (assuming annual interest payments).

Can this be input into TMV (time value of money) in the calculation, if so how?
Answer #1

Step 1: Price of the bond today
I/Y=12%
PMT=-7%*1000
FV=-1000
N=10
CPT PV=717.4888

Step 2: Price of the bond 1 year later
I/Y=8%
PMT=-7%*1000
FV=-1000
N=9
CPT PV=937.5311

Step 3: Return
=(937.5311+1000*7%)/717.4888-1
=40.425%